OpenStudy (anonymous):

What is the simplified fraction form of 0.44? 22/50 11/25 44/100 22/100

OpenStudy (anonymous):

0.44=44/100=22/50=11/25 so the answer is 11/25

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

welcome